# A Survey on Compositional Learning of AI Models: Theoretical and Experimental Practices

### v1

@@ 5 Theoretical Findings
Classical Neural Network. [95] provides a mathematical definition of compositionality for learning models and connects their expressively to computational complexity. They frame the existing well-known models, such as variations of RNN and CovNets, with the provided formal definition to explain properties related to their compositional generalization. [34] further investigates the RNN’s ability to generate natural language with a certain nesting depth. They claim that RNNs with optimal memory and $O(m\log k)$ hidden units can generate a natural language of well-nested brackets of $k$ types and $m$ bounded nesting depth. With the rise of LLMs, compositional generalization has become more critical recently. LLMs perform empirically well on many tasks due to their large-scale parameters and training data. However, the empirical performance measures are now less reliable, as the high performance on test data can not be interpreted as compositional generalization anymore. This issue is due to the nature of internet-scale training of LLMs and data contamination. Consequently, there is more urgency for theoretical studies to understand their limitations and measure their reliability in unobserved situations. However, [1] argued that studying the smaller models at the single neuron level potentially leads to a better understanding of the large/deep models’ learning behavior. They establish a connection between the Edge of Stability identified by the learning rate of the gradient descent approach for non-convex optimization and the emergent abilities in learning. This result remains limited to the scope of a single neuron and has not been extended to large models yet.
@@
Transformers. To establish the limitations of LLMs, it is essential to investigate the limitations of transformers, that is their underlying architectural component. In this work [73], the authors assume a specific transformer type, suggesting that their arithmetic precision is logarithmic in the number of input tokens. Based on this assumption, they demonstrate that transformers cannot accurately solve linear equalities or check membership in an arbitrary context-free grammar with empty productions. Additionally, transformer precision has been studied before in [15]. They claim that standard transformers have limited precision, implying that transformers cannot handle an infinite input length. Another notable theoretical investigation focuses on the activation functions to explain the limitation of the transformer [92, 15, 30]. [30] analyzes both hard-attention and soft-attention transformers. For hard attention, they prove that the transformer ignores most of the input information diagnosed by the specific modifications applied to the input. According to their analysis, it will be impossible for transformers with hard attention to solve problems that require processing the entire input, such as PARITY and logical formula problems. However, this conclusion contradicts older papers that state transformers are Turing complete [92, 91]. To prove Turing completeness, they utilize the strong assumption that the entire input information is accessible using hard attention. This leads to a different conclusion, stating that the transformer is capable of computing and accessing the entire internal dense representation. When using soft attention, they illustrate that it struggles with solving long input. They show that the influence of input on output substantially drops as the input gets longer. This is a similar conclusion, using a different approach analyzed in [15].
@@
Despite proving weaknesses of the transformer, [30] claims that the transformer has the potential to completely solve small input tasks. Two recent works also support this claim. The first work provides proof utilizing computation graphs and a theoretical study of error propagation in transformers. They claim that the auto-regressive transformer’s error reduces as the size of the input decreases [17]. Moreover, they show that transformers typically reduce problems into multi-step compositional problems to solve larger tasks. The second work supports the mentioned claim based on the study of sub-sequential finite-state transducers (SFSTs) [115]. They generate a set of random SFSTs following Montague’s Compositionality theorem to discover the coverage limitation. This limitation is inversely related to the size of the dataset and significantly impacts the probability of successful generalization of a model.
@@
Large Language Models. In addition to inconclusive theoretical studies on transformer limitations, there is also a discussion on large language models. The most noteworthy study is on the emerging abilities and capabilities that are claimed to be unique in the large models. The emerging abilities extend to the compositionality of the models. Although there are several experiments indicating these emerging capabilities, at least two papers disclaim their existence. The first group provides a theoretical proof based on a mathematical framework. They illustrate that the emerging ability appears due to the selected evaluation metrics that are nonlinear and discontinuous [99]. They show as an artifact of the evaluation metrics, even simple models such as CNNs can show emerging abilities. Therefore, they conclude that emerging abilities may not be a fundamental property of the large models. Moreover,  [68] provides an extensive empirical study with 1000 experiments on 22 tasks with different LLMs. However, they do not find any strong evidence of how the abilities emerge, given the inconsistency in some of the results and the unpredictability of emerging abilities. They associate the performance with in-context learning techniques, memorization, and data contamination. However, there is a recent publication that presents a positive theoretical analysis of reasoning capabilities by studying the chain of thought(CoT) [116], which draws a different conclusion. They argue that the log-precision transformer potentially possesses the ability to perform fundamental operations such as multiplication and a look-up table. Consequently, it can solve linear equations and other reasoning problems if it stores the entire input information. However, the architecture alone struggles with storing the entire input, as observed in [15, 73]. They show that the model addresses this limitation by repeatedly referring back to the input by enabling CoT [18]. Therefore, LLMs can overcome the weakness of the transformer in solving mathematical reasoning with the right amount of CoT examples.
@@
Model Type Theoretical Analysis Experimental Basic Neural Models RNN [34] [7] CNN [39] LSTM [104] Transformer-based (Customized) Transformers  [30][92],  [15] [83] LLM [17] [99] Neuro-symbolic Neural Modular Network  [57] Other Models  [29]
@@
Table 2: Summary of computational models with compositional learning ability from the theoretical perspective and an eaxmple from the experimental perspective.
@@ 6 Discussion and Future Direction
There has been a large amount of research on the compositional learning ability of humans from a cognitive perspective [19, 43]. Researchers in linguistics and formal languages have formalized the notion of compositionality since languages have inherent compositional structure [10, 11]. However, from the AI and machine learning perspective, ideas are borrowed from both cognitive and linguistics, and computational tasks and models are designed focusing on narrow aspects of compositionality [39]. Our investigation of AI models indicates several challenges regarding the designed tasks, benchmarks, and theoretical frameworks that make the evaluation of computational models problematic.
@@
Synthetic and Unrealistic Evaluations. One issue in current evaluations is that controlled and clean tests of compositonality are mostly synthesized [117, 97]. Even in rare cases that claim to work with realistic data [49], synthesized questions are used to query knowledge graphs. However, more recent studies on language models’ evaluation of compositionality focus on more challenging problems such as multi-hop question answering [89, 67, 82, 76] as well as complex puzzles with combinatorial search solutions or compositional mathematical reasoning [17].
@@
LLM Evaluation Challenge. The second challenge that mostly applies to LLMs is data contamination. Though the recent research compares language models to the specialized architectures and indicates their outperformance in compositional tasks  [21], this result does not necessarily mean these models have better generalizations in recognizing unobserved compositions [89]. A major issue with these evaluations on realistic data is the difficulty in disentangling the compositional reasoning from the data contamination and memorization. The generalization abilities can be an artifact of observing more complex data in larger contexts as well as evaluation metrics as has been pointed out in [99].
@@
Inconsistent Theoretical Methodology. Given the difficulty in obtaining conclusive empirical studies, theoretical understandings become even more important, nowadays. However, the lack of a well-established and practically informative theoretical framework for the investigation of the limitations and capabilities of LLMs has been a challenge to a deep understanding of their generalizability. According to our studies on the theoretical explanation of transformers, their compositionality is still under discussion. Some results illustrate that transformers possess compositional generalizability based on their ability to solve complex tasks based on smaller subtasks [18, 17]. However, some other results based on a different evaluation methodology suggest that the emergence of such abilities, including compositional learning, is unpredictable [68, 99]. Many empirical results confirm the latter view, for example, transformers still have serious limitations in comprehending large inputs [30, 15]. Despite these discussions, the number of studies on the theoretical analysis of transformer-based models is quite limited, especially the variety of procedures to prove the possibility and limitations of the models. Therefore, the compositional capabilities of the current SOTA models need more attention from the research community. This research direction will help towards more conclusive results on the limitations of the model’s generalizability, including the compositional generalization.
@@
Cognitive Motivation. The fundamental capabilities of current AI models have been debated and criticized by scientists in cognitive science and psychology [6, 70]. Despite giant leaps of performance progress in modern AI, there are distinct differences between these machines and human intelligence. Evaluating different models reveals that they often rely simply on pattern recognition [24, 17], instead of a holistic understanding of a problem grounded in reality and situation. Understanding human intelligence from Cognitive Science literature suggests that we must move beyond current engineering trends to build causal models of the world that support knowledge and understanding. The key ingredients of such human-like rich and efficient learning are compositionality and learning-to-learn [60].
@@ Acknowledgements
This project is partially supported by the Office of Naval Research (ONR) grant N00014-23-1-2417. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the Office of Naval Research. We would like to thank Dr. Tim Klinger (tklinger@us.ibm.com) for reviewing this manuscript and providing insightful feedback.

### v2

@@ 5.1 Classical Neural Network
Ram et al. (2023) provides a mathematical definition of compositionality for learning models and connects their expressively to computational complexity. They frame the existing well-known models, such as variations of RNN and CovNets, with the provided formal definition to explain properties related to their compositional generalization. Hewitt et al. (2020) further investigates the RNN’s ability to generate natural language with a certain nesting depth. They claim that RNNs with optimal memory and $O(m\log k)$ hidden units can generate a natural language of well-nested brackets of $k$ types and $m$ bounded nesting depth. With the rise of LLMs, compositional generalization has recently become more critical. Due to their large-scale parameters and training data, LLMs perform empirically well on many tasks. However, the empirical performance measures are now less reliable, as the high performance on test data can not be interpreted as compositional generalization anymore. This issue is due to the nature of internet-scale training of LLMs and data contamination. Consequently, there is more urgency for theoretical studies to understand their limitations and measure their reliability in unobserved situations. However, Ahn et al. (2023) argued that studying the smaller models at the single neuron level potentially leads to a better understanding of the large/deep models’ learning behavior, which is related to explaining the Systematicity of the model. They also establish a connection between the Edge of Stability identified by the learning rate of the gradient descent approach for non-convex optimization and the emergent abilities in learning. This result remains limited to the scope of a single neuron and has not yet been extended to large models.
5.2 Transformers
@@
To define the limitations of LLMs, it is essential to investigate the limitations of transformers and their underlying architectural component. In this work (Merrill & Sabharwal, 2023), the authors assume a specific transformer type, suggesting that their arithmetic precision is logarithmic in the number of input tokens. Based on this assumption, they demonstrate that transformers cannot accurately solve linear equalities or check membership in an arbitrary context-free grammar with empty productions. The studies of transformer precision have been explored before in Dehghani et al. (2019). They claim that standard transformers have limited precision, implying that they cannot handle an infinite input length. This conclusion indicates the limitations in of the compositionally of the transformers in terms of the Productivity aspect. Another notable theoretical investigations focus on the activation functions to explain the limitation of the transformer (Dehghani et al., 2019; Hahn, 2020). Hahn (2020) analyzes both hard-attention and soft-attention transformers. For hard attention, they prove that the transformer ignores most of the input information diagnosed by the specific modifications applied to the input. According to their analysis, transformers with hard attention will be unable to solve problems that require processing the entire input, such as PARITY and logical formula problems. However, this conclusion contradicts older papers that state transformers are Turing complete (Pérez et al., 2021). They utilize the strong assumption that all input information is accessible using hard attention to prove Turing completeness. This leads to a different conclusion, stating that the transformer can compute and access the entire internal dense representation. Hahn (2020) also investigate the model’s behavior with soft attention. They illustrate that Transformers struggle with solving long input by demonstrating the influence of input on output substantially drops as the input gets longer. This is a similar conclusion, using a different approach analyzed in an older study of (Dehghani et al., 2019). Based on these analyses, they further confirm the lack of productivity aspect of transformer architecture caused by limited training data lengths.
@@
Despite proving weaknesses of the transformer, Hahn (2020) claims that the transformer has the potential to solve small input tasks completely. Two recent works also support this claim. The first work provides proof utilizing computation graphs and a theoretical study of error propagation in transformers. They claim that the auto-regressive transformer’s error reduces as the size of the input decreases (Dziri et al., 2023). Moreover, they show that transformers reduce problems into multi-step compositional problems to solve larger tasks, which is strongly related to the Novel Composition of the compositional aspects. The second work supports the mentioned claim based on the study of sub-sequential finite-state transducers (SFSTs) (Valvoda et al., 2022). They generate a set of random SFSTs following Montague’s Compositionality theorem to discover the coverage limitation. This limitation is inversely related to the size of the dataset and significantly impacts the probability of a model’s successful generalization.
5.3 Large Language Models
@@
In addition to inconclusive theoretical studies on transformer limitations, there are controversial results on large language models. The most noteworthy study is on the emerging abilities and capabilities claimed to be unique in the large models. The emerging abilities relate to the generalization to new and complex tasks in LLMs. This kind of ability is also a feature of models’ compositional learning ability, allowing them to perform in novel compositional situations (Yu et al., 2024). Multiple works have shown the existence of emergent abilities of LLMs (Wei et al., 2022). The recent work of Arora & Goyal (2023) provides a mathematical framework for identifying complex skills in language models. They use the LLM Scaling Rule to argue that emergent skills are the results of reducing excessive loss. This excessive loss enables the model to learn how to utilize and combine skills from downstream tasks during training. Their claims are based on the assumption that language inherently contains a random mix of complex skills. Although several experiments reveal these emerging capabilities, at least two papers disclaim their existence. The first group provides a theoretical proof based on a mathematical framework. They illustrate that the emerging ability appears due to the selected evaluation metrics that are nonlinear and discontinuous (Schaeffer et al., 2023). They show as an artifact of the evaluation metrics, even simple models such as CNNs can show emerging abilities. Therefore, they conclude that emerging abilities may not be a fundamental property of the large models. Moreover,  Lu et al. (2024) provides an extensive empirical study with 1000 experiments on 22 tasks with different LLMs. However, given the inconsistency in some results and the unpredictability of emerging abilities, they do not find any strong evidence of how they emerge. They associate the performance with in-context learning techniques, memorization, and data contamination. However, a recent study presents a positive theoretical analysis of reasoning capabilities by investigating the chain of thought (CoT) (Wei et al., 2024) and draws a different conclusion. They argue that the log-precision transformer can perform fundamental operations such as multiplication and a look-up table. Consequently, it can solve linear equations and other reasoning problems if it stores all the input information. However, the architecture alone struggles with storing the entire input, as observed in Dehghani et al. (2019); Merrill & Sabharwal (2023). They show that the model addresses this limitation by repeatedly referring to the input by enabling CoT (Feng et al., 2023). Therefore, with the right number of CoT examples, LLMs can overcome the transformer’s weakness in solving mathematical reasoning. This is aligned with previous empirical results of in-context learning discussed in Section 4.2.
@@
Model Type Theoretical Analysis Empirical Basic Neural Models RNN Hewitt et al. 2020 Bowman et al. 2015 CNN Hupkes et al. 2020 LSTM Siegelmann & Sontag 1995 Transformer-based (Customized) Transformers  Hahn 2020Pérez et al. 2019,  Dehghani et al. 2019 Ontanon et al. 2022 LLM Dziri et al. 2023 Schaeffer et al. 2023 Neurosymbolic Neural Modular Network  Kuo et al. 2020 Other Models  Gupta & Kembhavi 2023
@@
Table 2: Summary of computational models with compositional learning ability from the theoretical perspective and an example from the experimental perspective.
@@ 6 Discussion and Future Direction
There has been a large amount of research on the compositional learning ability of humans from a cognitive perspective (Fodor & Pylyshyn, 1988; Ito et al., 2022). Researchers in linguists and formal languages have formalized the notion of compositionality since languages have inherent compositional structure (Chomsky, 1965; 2002). However, from the AI and machine learning perspective, ideas are borrowed from both cognitive and linguistics, and computational tasks and models are designed focusing on narrow aspects of compositionality (Hupkes et al., 2020). Our investigation of AI models indicates several challenges regarding the designed tasks, benchmarks, and theoretical frameworks that make the evaluation of computational models problematic.
Figure 1 shows the connections between the main topics identified and discussed in this survey. Among the five identified types of compositional learning facets, only systematicity and productivity are well-researched and have clear connections to evaluation benchmarks. Given that these are the five main metrics of compositionality, we should aim to expand our evaluation capabilities by developing tests for the other three as well. Empirical evaluations are comparatively more well-studied compared to theoretical analyses. Theoretical evaluations are either lacking or do not follow a consistent methodology. There is a lack of connection between the theoretical methods and the cognitive aspects, making these results hard to use to guide better architectural design. For Models, different types of architectures have been designed. However, evaluating LLMs and making fundamental design decisions for compositional generalization present new challenges. We outline some of these challenges in detail below.
Less Explored Facets of Compositionality. Only Systematicity and Productivity have been well-researched and have established connections to evaluation benchmarks. While the other three were introduced as fundamental types of compositionality, they have received less attention, as they appear to be less commonly occurring aspects of compositionality. However, in the era of LLMs and the emergent in-context learning, Substitutivity and Localism are potential bottlenecks in the performance of LLMs for attending the appropriate context for solving problems. Moreover, Overgeneralization can be associated with hallucination and generating unfounded and incorrect information by making up new unrealistic abstractions. While hallucination is a broader concept than overgeneralization, this compositional learning facet can highlight an important aspect to be addressed to prevent hallucination (Huang et al., 2023). Therefore, directing attention to the other three types of measures can help establish new formal evaluation benchmarks, contributing to the development of more robust systems and addressing the challenges of the LLMs.
@@
Synthetic and Unrealistic Evaluations. One issue in current evaluations is that controlled and clean tests of compositonality are mostly synthesized (Wu et al., 2021; Ruis et al., 2020). This is evidenced by our examples in Section 3. Even in rare cases that claim to work with realistic data (Keysers et al., 2020), synthesized questions are used to query knowledge graphs. However, more recent studies on language models’ evaluation of compositionality focus on more challenging problems such as multi-hop question answering (Press et al., 2023; Liu et al., 2022b; Okawa et al., 2023; Mirzaee et al., 2021) as well as complex puzzles with combinatorial search solutions or compositional mathematical reasoning (Dziri et al., 2023). Although the benchmarks are designed for evaluating specific tasks, the reliance on mostly synthesized data risks the effectiveness of generalization to real-world data, which is often more complex.
@@
Misalignment of Performance and Compositional Learning (LLM Evaluation Challenge). The second challenge that mostly applies to LLMs is data contamination. Though the recent research compares language models to the specialized architectures and indicates their outperformance in compositional tasks (Furrer et al., 2021), this result does not necessarily mean these models have better generalizations in recognizing unobserved compositions (Press et al., 2023). A major issue with these evaluations on realistic data is the difficulty in disentangling the compositional reasoning from the data contamination (Sainz et al., 2023) and memorization. However, very little recent work has been done to address this issue (Yu et al., 2024). Based on our investigation of theoretical studies in Section 5, we identify that the generalization abilities can be an artifact of observing more complex data in larger contexts as well as evaluation metrics as has been pointed out in Schaeffer et al. (2023) and is a pertinent issue that needs to be addressed in new evaluation benchmarks.
@@
Inconsistent Theoretical Methodology. Given the difficulty in obtaining conclusive empirical studies, theoretical understandings become even more important, nowadays. However, the lack of a well-established and practically informative theoretical framework for investigating the limitations and capabilities of LLMs has been a challenge to a deep understanding of their generalizability. According to our studies on the theoretical explanation of transformers in Section 5, their compositionality is still under discussion. Some results illustrate that transformers possess compositional generalizability based on their ability to solve complex tasks based on smaller subtasks (Feng et al., 2023; Dziri et al., 2023). However, other results based on a different evaluation methodology suggest that the emergence of such abilities, including compositional learning, is associated with the user’s choice of evaluation metrics (Schaeffer et al., 2023). Similar lines of study deny the emergence of intelligence and relate the new abilities to in-context learning methods, models memory, and linguistic knowledge (Lu et al., 2024). Many research results confirm the limitations in the generalizability of LLMs. For example, the building blocks of these models, i.e. transformers, still have severe limitations in comprehending large inputs (Hahn, 2020; Dehghani et al., 2019). Despite these controversial discussions, there are only relatively few studies on the theoretical analysis of transformer-based models. The methodological frameworks for examining the LLMs’ capabilities are not standardized. Therefore, the compositional capabilities of the current SOTA models require more attention from the research community. This research direction will help towards more consistent and conclusive results on the limitations of the model’s generalizability, specifically the compositional generalization.
@@
Cognitive Motivation. The fundamental capabilities of current AI models have been debated and criticized by scientists in cognitive science and psychology (Bender & Koller, 2020; Marcus, 2018). Despite giant leaps of performance progress in modern AI, there are distinct differences between these machines and human intelligence. From our exploration of cognitive aspects of compositional learning in Sections 1 and 2, we observe that cognitive foundations do not have strong ties to model building yet. Evaluating different models reveals that they often rely simply on pattern recognition (Geirhos et al., 2020; Dziri et al., 2023), instead of a holistic understanding of a problem grounded in reality and situation, as was seen in Section 5. Understanding human intelligence from Cognitive Science literature suggests that we must move beyond current engineering trends to build causal models of the world that support knowledge and understanding. The key ingredients of such human-like rich and efficient learning are compositionality and learning-to-learn (Lake et al., 2017).
7 Limitations
Despite the comprehensive nature of the survey and our efforts to cover and connect most research relating to compositional learning, we would like to acknowledge some limitations. The scope of this survey covers a broad spectrum of topics and tries to capture both theoretical and experimental frameworks, but there might be some relevant papers that are missed. Compositional learning is an interdisciplinary topic across Computer Science, Linguistics, Cognitive Science, etc. Although we have included insights and connections from across these fields, our work has a more in-depth focus on Computer Science literature, especially Natural language processing. While we tried to provide the overall picture of the related research and build a coherent story, we might not capture the detailed nuances of each definition and application.
@@ Acknowledgements
This project is partially supported by the Office of Naval Research (ONR) grant N00014-23-1-2417. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the Office of Naval Research. We thank Dr. Tim Klinger and the anonymous reviewers for their constructive feedback, which greatly helped us improve this manuscript.